OpenStudy (anonymous):

On a manufacturer's assembly line, 3% of its merchandise comes off defective. If the company has 354 defective items, how many total items came off the assembly line?

OpenStudy (lucaz):

3% of n = 354 \[\large{\frac{3}{100}(n)=354\\ 3n=354\times 100\\ n=\frac{35400}{3}=11800 }\]